OpenStudy (anonymous):

What are the dimensions of a rectangle that has a perimeter of 56 units and an area of 96 square units? a. 4 by 24 c. 14 by 14 b. 2 by 48 d. 3 by 32

OpenStudy (hoblos):

suppose the dimensions of the rectangle are length l and width w perimeter = 2l + 2w area = l*w thus 2l + 2w = 56 l*w = 96 we have a system of two equations, can you solve it ?

OpenStudy (anonymous):

a. 4 by 24 c. 14 by 14 b. 2 by 48 d. 3 by 32